Peter travels from his place to his uncle house a distance of 30km.He cycles for two thirds of the journey before walking for the rest of the journey. If his cycling speed is 10km/hr more than the walking speed and the whole journey took him 3hrs 20 minutes, calculate his cycling speed.


1
Expert's answer
2021-02-24T07:50:26-0500

"3\ hr\ 20\ min=\dfrac{10}{3}\ hr"

Let "x" km/hr be his walking speed. Then his cycling speed is "(x+10)" km/hr.


"\dfrac{\dfrac{2}{3}\cdot30}{x+20}+\dfrac{\dfrac{1}{3}\cdot30}{x}=\dfrac{10}{3}"

"6x+3(x+10)=x(x+10)"

"x^2+x-30=0"

"(x+6)(x-5)=0"

"x_1=-6, x_2=5"

Since "x>0," we take "x=5."

Then


"x+10=5+10=15"

His cycling speed is 15 km/hr.